OpenStudy (anonymous):

I really need help on this math problem. I was wondering if anyone can show me what I'm doing wrong, the problem is, 7x-7y=3 solve for y ..I keep getting 3+x but the answer is y=x-3/7

OpenStudy (anonymous):

7x - 7y = 3 -7y = -7x + 3 //divide both sides by -7 y = x - 3/7

OpenStudy (anonymous):

but i dont get it, wouldnt there be a 7 under the x, also?

OpenStudy (anonymous):